OpenStudy (anonymous):

OK so if i buy 12 oz of weed and it is $25 for one oz how much is the cost and if i only have $500 how many more oz. can i buy?

OpenStudy (anonymous):

12oz * $25 = 300 500 - 300 = 200/25 = 8